### v2

# Possibilities for methanogenic and acetogenic life in molecular cloud

Abstract
According to panspermia, life on Earth may have originated from life forms transported through space from elsewhere. These life forms could have passed through molecular clouds, where the process of methanogenesis could have provided enough energy to sustain living organisms. In this study, we have calculated the Gibbs free energy released from synthesizing hydrocarbons for methanogenic (acetogenic) life in a molecular cloud, with methane (acetic acid) as the final metabolic product. Our calculations demonstrate that the chemical reactions during methanogenesis can release enough free energy to support living organisms. The methanogenic life may have served as the ancestor of life on Earth, and there is some preliminary evidence from various molecular biology studies to support this idea. Furthermore, we propose a potential distinguishing signal to test our model.
I Introduction
In a recent paper, the authors found that the last universal common ancestor (LUCA) appeared 4.2 billion years ago which is just 400 million years after the formation of Earth LUCA . This may be a hint of panspermia. According to panspermia Panspermia , life on Earth could have descended from forms of life transported through space. Such life forms could have passed through and survived for some time within molecular clouds. Hydrogen molecules maintain a liquid state within the temperature range of 13.99 K to 20.27 K, which coincidentally coincides with the typical temperature of molecular clouds. Although it is possible to assume the liquid hydrogen environment of molecular cloud life, more serious problems arise correspondingly. So we won’t discuss such a possibility here. Such extremely low temperatures environment may have unexpected benefits. In Ref   Chirality , the author argues that the ultra-low temperature environment of molecular clouds may be the reason for the chiral polymer chain of biological molecules. For more information about the effects of low temperatures on the origin of life, see Ref. lowt1 ; lowt2 .
A crucial question arises concerning the energy acquisition strategies employed by the cloud lifeforms. The author has previously suggested a bioenergetic mechanism driven by cosmic rays cosmic-ray-bioenergetics , which relies on the ionization of hydrogen molecules as its primary energy source. However, it is worth exploring alternative energy acquisition pathways that may support life in these extreme environments.
Metabolic activity and reproduction of biological systems require energy transformations through respiration processes, such as aerobic and anaerobic respiration. Aerobic and anaerobic respiration are all redox reactions of biological fuels in the presence of an inorganic electron acceptor. The difference between them is whether the electron acceptor is oxygen. Such redox reactions release Gibb Free energy and produce large amounts of energy stored in ATP.
On the Earth, methanogenic bacteria reduce $\rm CO_{2}$ to $\rm CH_{4}$ and release Gibbs free energy needed for life through methanogenesis, which is anaerobic respiration with methane as the final product of metabolism and only found in the domain Archaea. C.P. McKay and H.D. Smith explored the biochemical reactions of hypothetical methane-based life on Titan, with reactants including $\rm C_{2}H_{2}$, $\rm C_{2}H_{6}$, and organic haze McKay2005 . Given the presence of similar compounds in molecular clouds, it is plausible to consider methanogens thriving in these cosmic regions as well. Here, we will discuss such probability and calculate the release of free energy for methanogenic life in the environment of molecular clouds. Since acetogenic life has the same electron donor and acceptor through the Wood-Ljungdahl pathway on Earth, the energy release of such biochemical processes in molecular clouds is also discussed. Then, we discuss the relationship between molecular cloud life and LUCA.
This paper is organized as follows: In Sec.2, we explore the possible biochemical reactions of methanogenic and acetogenic life in molecular clouds and calculate the released Gibbs free energy. A possible distinguishing signal is also presented in this section. The relation between such methanogenic and acetogenic life in molecular clouds and LUCA are discussed in Sec 3.. The conclusions are summarized in the final section.
II Possibilities for methanogenic life in molecular cloud
II.1 The biochemical reactions of methanogenic and acetogenic life in molecular clouds
Methanogens on Earth maintain their activity through methanogenesis, a form of anaerobic respiration. Carbon dioxide is the terminal electron acceptor in methanogenesis through the following chemical reaction
$\rm{CO_{2}+4H_{2}\rightarrow CH_{4}+2H_{2}O}.$ (1)
With the same electron acceptor and donor, acetogenic processes are also possible through the following reaction
$\rm{2CO_{2}+4H_{2}\rightarrow CH_{3}COOH+2H_{2}O}.$ (2)
As the $\rm CO_{2}$ in molecular clouds is more abundant in the solid phase than in the gas phase Minh1988 ; Gerakines ; Whittet , solid $\rm CO_{2}$ provides abundant carbon sources and energy for metabolic activities. Then there may be life forms gathered around these solid forms of carbon dioxide.
The most abundant molecule in molecular clouds is H2 which accounts for about 70%. It is also the main collision target of other molecules therein. The second most abundant molecule in the molecular cloud is $\rm CO$, and its hydrogenation reaction, as in Eq. 3 may be the energy source of methanogen-like creatures.
$\rm{CO+3H_{2}\rightarrow CH_{4}+H_{2}O}.$ (3)
In addition, we also considered the chemical reaction given in Ref. McKay2005 , which is
$\rm{C_{2}H_{2}+3H_{2}\rightarrow 2CH_{4}}.$ (4)
The Gibbs free energy, $\Delta G$, released from these two reactions is computed for molecular cloud conditions using the method described in Ref. Kral1998 , which is
$\Delta G=\Delta H-T\Delta S+RT{\rm ln}(Q),$ (5)
where $\Delta H$ ($\Delta S$) is the difference in heats of formation (entropy) of product and reactant under standard conditions, which is shown in Tab. 1, $T$ is the temperature, and $R$ is the universal gas constant. $Q$ is the ratio of the activities between products and reactants raised to the power of its multiplying constant in the chemical reaction equation. Following Ref. McKay2005 , the chemical activities of the gas molecules we used in this draft are approximately equal to their partial pressure in molecular clouds. The activity of solid carbon dioxide is unity.
Table 1: Heats of formation and entropy at standard conditions ($\rm 25^{\circ}$, 1 bar)
                         Molecule $H({\rm kJ~{}mol^{-1}})$ $S({\rm J/mol~{}K})$ Abundance
                         $\rm H_{2}(g)$ 0 130.68Hydrogen $1.1\times 10^{4}~{}n_{\rm CO}$rerking_1982
                         $\rm CH_{4}(g)$74.6Methane 186.3Methane ${\rm 10^{-3}~{}}n_{\rm CO}$acy_1991
                         $\rm C_{2}H_{2}(g)$ 227.400Methane 200.927Methane ${\rm 3\times 10^{-4}\sim 10^{-3}}~{}n_{\rm CO}$acy_1989
                         $\rm H_{2}O(g,gas)$241.83 NIST 188.84 NIST ${\rm 10^{-9}\sim 10^{-7}}~{}n_{\rm H_{2}}$(for ortho-$\rm H_{2}O$)Snell_1999
                         $\rm CO(g)$110.53 NIST 197.66 NIST $n_{\rm CO}$
                         $\rm CO_{2}(g,solid)$427.4 51.07Giauque -
                         $\rm CH_{3}COOH$433 282.84Weltner ${\rm 3.4\times 10^{-10}}~{}n_{\rm H_{2}}$ehringer
The typical temperature of a molecular cloud is about $\rm 10\sim 20~{}K$, but temperatures of ”warm clouds” are about 20-60 K and molecular clouds with H II regions can reach 100 K temperature . So, we set $\rm 10~{}K<T<100~{}K$ in this draft. The partial pressure of $\rm H_{2}$, $\rm CH_{4}$, $\rm C_{2}H_{2}$, $\rm H_{2}O$, $\rm CO$ and $\rm CH_{3}COOH$ we used here are 70%, $6\times 10^{-8}$, $3\times 10^{-8}$, $7\times 10^{-9}$, $6.36\times 10^{-5}$ and $2.38\times 10^{-10}$, respectively.
Figure 1: The Gibbs free energy released from the synthesis of hydrocarbons.
In Fig. 1, we show the Gibbs free energy released from these reactions in the molecular cloud environment. The change of Gibbs free energy is about $\rm-60\sim-370~{}kJ/mol$. A negative value of Gibbs free energy is a spontaneous chemical reaction that releases energy. The minimum energy to sustain the survival of methanogen on Earth is about $\rm 42kJ/mol$ Kral1998 . And the Gibbs free energies released by the reactions we considered here are energetically acceptable.
However, the density of acetylene in molecular clouds is too low to provide enough free energy. Although carbon monoxide is much denser, whether it can provide enough energy depends on the strength of metabolic activity at such low temperatures. Methanogenic and acetogenic life attached to solid carbon dioxide does not have the problem of insufficient raw materials. Solid carbon dioxide can provide sufficient energy and carbon sources for metabolic activities. In addition, there may be other bioenergetics mechanisms as a supplement, such as the mechanism driven by cosmic ray ionization cosmic-ray-bioenergetics . Different bioenergetics mechanisms may work together. However, the processes discussed here have unique advantages in fixing carbon by converting carbon elements to organic molecules for life in molecular clouds using methane as an intermediate material.
II.2 A possible distinguishing signal
The consumption and production of carbon compounds through metabolic activities may significantly affect the distribution of these molecules. It could potentially serve as a trace signal of life within molecular clouds. According to this hypothesis, molecular cloud life uses solid carbon dioxide (carbon monoxide or $\rm C_{2}H_{2}$) as a stable carbon source, which causes significant methane or acetic acid production cosrrespondingly. As a result, it can be predicted that the distribution of methane (or acetic acid) in molecular clouds would mirror that of solid carbon dioxide. As the protoplanetary disk formed in molecular cloud and the life forms therein would be concentrated correspondingly, the consistency effect of the molecular distribution mentioned above will be more significant. The same conclusion also applies to the distribution of carbon monoxide and $\rm C_{2}H_{2}$. If such distribution consistency does not exist, it would be a great challenge for our model.
In fact, some studies have shown that carbon dioxide and methane are distributed in the space region of two protostars: IRAS 16253-2429 Beuther2023 and IRAS 23385+6053 Narang2023 . If future astronomical observations confirm this distribution consistency, it should prompt consideration of an astrochemical origin as the first possibility. After the ruling out of astrochemical origin, it could serve as a verification of the model.
III The relationship with LUCA
If there were methanogenic or acetogenic life in the pre-solar nebula and then fell on the early Earth, LUCA would be one type of methanogenic or acetogenic life. Since acetogenic bacteria are usually considered a type of bacteria rather than archaea, LUCA is more inclined towards a certain type of methanogenic life. The three-domain tree of life presented by ribosomal RNA ribosomal-RNA depicted that LUCA is the last common ancestor of archaea, eukaryotes and bacteria. In that framework, efforts to infer the gene content, hence the properties of LUCA, boiled down to identifying genes that were present in eukaryotes, archaea, and bacteria. Coincidentally, the primitive atmosphere of early Earth was rich in carbon monoxide and carbon dioxideKASTING . It ensures the continued survival of methanogenic (acetogenic) life on Earth. Some previous studies Pierce ; Rother have demonstrated that certain acetogenic and methanogenic Archaea utilized carbon monoxide as a carbon source during the early stages of Earth’s development.
The most important prediction of this model is that these molecular cloud life could be the ancestor of life on Earth. More generally, LUCA is one type of life form that uses $\rm H_{2}$ ($\rm CO$ or $\rm CO_{2}$) as an electron donor (acceptor) in the chemical reactions of generating energy. In general, this model provides a perfect origination of LUCA. There is already some tentative evidence about the probability that LUCA is one type of methanogenic or acetogenic life. Some researches indicate that early lifeforms invented methanogenesis in the cooler temperature zones that led to the rise of LUCA after acclimatization to the hydrothermal vent environment by identifying the minimum genome of LUCA mat . It has been reported that LUCA’s genes point to methanogenic and acetogenic roots by surveying nearly two thousand genomes of modern microbes Madeline . Despite a few challenging studies Berkemer , research on geological evidence and phylogenomic reconstructions supports this thesis Preiner ; Xavier .
In this model, the descendant of this ancient life would exist in all places in the solar system suitable for its survival, such as Europa and Titan. It should be noted that there is no suggestion here that life originated in molecular clouds. We believe that molecular cloud life may have derived from other planets in the Milky Way or planets in the solar’s predecessor star system nebula-relay .
IV Summary
This paper explores the possibility of methanogenic and acetogenic life in molecular clouds. The calculations demonstrate that the reaction of carbon monoxide, carbon dioxide, or acetylene with hydrogen molecules releases sufficient Gibbs free energy to ensure the survival of molecular cloud life. Solid $\rm CO_{2}$ may be the primary energy and carbon source because of its high local density. As the second most abundant molecule, the reaction of carbon monoxide is also a good option for providing energy. Additionally, we propose a potential distinguishing signal that could either support or disprove this model. We also consider the possibility that these methanogenic and acetogenic forms of life could have been LUCA’s interstellar ancestors. Even now, their descendants and fossils may still widespread in the solar system.
Acknowledgements.
We thank Dr. Yan Sun and Bing-Yu Su for their generous help in perfecting this article. This work is supported by the National Natural Science Foundation of China (Grants No. 12373002, 11773075) and the Youth Innovation Promotion Association of Chinese Academy of Sciences (Grant No. 2016288).